Find the sum of vector A and B
A=16i + 16j
B=16i - 32j
1
Expert's answer
2020-02-24T10:49:38-0500
"\bold{A+B}=(16\bold{i}+16\bold{j})+(16\bold{i}-32\bold{j})"

"\bold{A+B}=(16+16)\bold{i}+(16-32)\bold{j}"

"\bold{A+B}=32\bold{i}-16\bold{j}"